OpenStudy (anonymous):

For which equation is x = 9 a solution? A. 3x – 7 = –34 B. x – 9 = 1 C. x/9 - 8 = 9 D. 4x + 5 = 41

OpenStudy (ivettef365):

substitute x for 9, to see which equation is true, for example, lets take the first one 3x - 7 = -34 3(9) - 7 = 34 27 - 7 = 34 20 = 34 (not true) keep trying with the other ones.

OpenStudy (anonymous):

okay, thanks